# Ineq-Comp: Benchmarking Human-Intuitive Compositional Reasoning in Automated Theorem Proving on Inequalities

### v1

@@ 4 Probing and Mitigating Compositional Challenges
To better understand and potentially mitigate the compositional failures observed in Section 3, we conduct two ablation studies. First, we explore whether our curated benchmark becomes easier when models are provided with the formal proof of the seed problem as part of the input (Section 4.1). This tests whether models can leverage in-context demonstrations to generalize to the transformed variants. Building on this idea, we then construct synthetic data that explicitly encodes compositional structure and use it to fine-tune the theorem prover. This examines whether the lack of compositional ability can be remedied through supervised fine-tuning alone (Section 4.2).
@@ 4.1 In-context learning
Table 2 reports the pass@$N$ accuracy across a range of open-source and proprietary models on Ineq-AMGM, Ineq-Cauchy, and Ineq-MISC problems under different generation budgets. Despite having access to the seed solution, all models continue to struggle with the transformed variants. For example, Qwen2.5-Coder-32B-Instruct achieves only 40.0% accuracy on Type II AM-GM problems and 20.0% on Type II Cauchy–Schwarz problems at 128 attempts. DeepSeek-R1-Distill-Qwen-32B (with thinking) performs similarly, reaching just 16.0% on Type I AM-GM problems. Proprietary models such as GPT-4o and Claude 3.7 Sonnet show mixed results, further illustrating the challenge. These findings suggest that simply providing the seed proof is not enough to enable compositional generalization. Even modest transformations remain difficult, pointing to a fundamental limitation in their ability to transfer and reuse formal reasoning strategies.
Table 2: ICL-based ablation. Performance of various models on Ineq-Simp (Type I and Type II variants) under the in-context learning (ICL) setting. For each test instance, the model is provided with the full formal proof of the corresponding seed problem as part of the input prompt. This setup evaluates whether models can leverage prior solutions to solve transformed variants via compositional generalization. We report pass@$N$ accuracy (with standard deviation as subscript) across different generation budgets $N$. Despite access to a complete solution to the seed problem, all models—including proprietary ones—continue to struggle on both Type I and Type II variants, indicating that the compositional gap cannot be easily closed through in-context demonstrations alone.
@@ 4.2 Fine-tuning
We split the seed problems into two subsets: 25 AM-GM inequality seed problems as training set and 50 other problems as held-out set. To construct training data, we use our automated framework (Ineq-Mix) to generated 15k problems incorporating both algebraic and compositional transformations base on the AM-GM seed problems. Then we used in-context learning (ICL) to prompt Qwen2.5-Coder-32B-Instruct with seed problems and their verified Lean 4 proofs. We retained approximately 8k compilable outputs and fine-tuned Goedel-Prover-SFT on this data. See Figure 4 for an overview of the process and Appendix D for setup details.
@@ 5 Related Works
ATP has been a long-standing goal in AI (Wu, 2008). Modern theorem provers use tactic-based proof search and premise selection within these systems (e.g. Lean’s tactics in Mathlib environment), but they struggle with enormous search spaces and limited training data for complex theorems (Polu and Sutskever, 2020; Yang et al., 2023). More recently, more theorem provers came out based on large language models (Xin et al., 2024; Lin et al., 2025; Dong and Ma, 2025; Wang et al., 2025; Ren et al., 2025), and also possibly incorporated tree-search techniques (Wu et al., 2024; Li et al., 2024; Xin et al., 2025).
@@ Appendix B More Details on Ineq-Mix
Positive RHS: if rhs $\text{if rhs }g(x_{1},\dots,x_{n})\text{ is guaranteed to be positive.}$.
@@
Table 3: (A list of composition methods for two inequality problems.) Given two inequality problems $P_{1},P_{2}$, where $P_{1}$ has variables $X=\{x_{1},\dots,x_{m}\}$, with the statement: given condition on the variables $C_{1}(X)$, the following inequality holds $f_{1}(X)\geq g_{1}(X)$, and $P_{2}$ has variables $X$, with the statement: given condition on the variables $C_{2}(X)$, the following inequality holds $f_{2}(X)\geq g_{2}(X)$, the following list defines several composition operations to generate a new inequality problem. If problems $P_{1}$ and $P_{2}$ have different set of variables, we can always lift the problems and make sure they are defined on the same variable space. Following are the composition methods we considered to combine $P_{1}$ and $P_{2}$ together. If $g_{1}(X)$ and $g_{2}(X)$ are both greater than 00, there are more composition methods. We also make sure that $C_{1}(X)\cap C_{2}(X)$ is non-empty after composition.
@@
Now we show the details of how to compose two inequality problems together or add algebraic transformation to create new inequality problems in Ineq-Mix. Recalled that as mentioned in Section 2.2, Ineq-Mix consists of transformation/composition rules from three categories: (1) compositional transformations, where (different) inequalities are combined through operations like addition, multiplication, or taking the maximum or minimum of both sides; (2) variable-level algebraic transformations, where we replace variables with other algebraic expressions; and (3) problem-level algebraic transformations, where we apply a monotone function to both sides of an inequality, such as taking exponentials or logarithms. We now introduce them by category.
Compositional transformations Composition transformations denote the rules that given two inequality problems, we compose them together to generate a new inequality, where the two inequality problems serve as subparts for the new inequality problem. Table 3 shows the full list of composition rules. Note that the rules are classified into two categories to make sure mathematical correctness: the rules require the right-hand side to be positive, and the rules without this requirement. Also note that if the two problems both has positive right-hand side in the statement, then after the compositions defined in Table 3, the new problem also has a positive right-hand side.
@@
Variable-level algebraic transformations Variable-level algebraic transformations denote the rules that apply algebraic transformations on the variables of a given inequality problem. For example, replacing $x_{i}$ in the original inequality problem into $x_{i}^{2}$, for all $x_{i}\in X$ where $X=\{x_{1},\dots,x_{m}\}$ denotes the variable set. Table 4 summarizes the variable-level algebraic transformations used for Ineq-Mix. Similar to compositional transformations, there are also two categories for variable-level algebraic transformations. The first category includes the transformations that are “one-to-one” mapping under the basic assumption $X\succ 0$, like taking the square root of all variables, and squaring all variables. Since these transformations are one-to-one, as long as the original problem is feasible, the new problem that applies these rules is also feasible. We also include some transformation rules that are not one-to-one under the basic assumption $X\succ 0$ (but are also elegant transformations). However to ensure the correctness, these rules can only be applied to problems where there is only the basic assumption $X\succ 0$ and no additional assumptions.
@@
Table 5 summarizes all the problem-level algebraic transformations used in Ineq-Mix.
@@ Appendix C Detailed Model Information
The full list of models used in our experiments, including the version (for API models) and Huggingface-link, is shown in Table 6.
Table 6: List of models tested in Section 3 or Section 4. We use “Kimina-Distill-7B” to denote “Kimina-Prover-Preview-Distill-7B” in the paper.
@@ D.1 Lean 4 version
For evaluation on models except InternLM2.5-StepProver, we stick to the Lean 4 and Mathlib version used in Xin et al. [2024], since Goedel-Prover-SFT [Lin et al., 2025] and STP [Dong and Ma, 2025] are trained based on DeepSeek-Prover-V1.5-RL [Xin et al., 2024]. Besides, we also nearly recover the reported performance of Kimina-Distill-7B [Wang et al., 2025] and DeepSeek-Prover-V2 [Ren et al., 2025] on MiniF2F benchmark using Lean 4 and Mathlib version in Xin et al. [2024]. We also observe decent performance of DeepSeek-R1-Distill-Qwen-32B on MiniF2F.
[…]
@@ D.4 Detailed results for fine-tuning Goedel-Prover-SFT
Figure 4: Pipeline for evaluating whether composition can be learned through naive fine-tuning. We split the seed problems into ${\mathcal{P}}_{\text{train}}$, which contains 25 problems utilizing AMGM, and ${\mathcal{P}}_{\text{held-out}}$, which contains the remaining 50 problems. The pipeline consists of three steps: (1) generate data by prompting LLM given the proof of seed problems in context and keep only the proofs that pass Lean compilation; (2) fine-tune Goedel-Prover-SFT; (3) evaluate type I and II problems and Ineq-Mix on ${\mathcal{P}}_{\text{held-out}}$
Please refer to Figure 4 for an illustration for our fine-tuning experiments pipeline. To generate the fine-tuning data ${\mathcal{D}}_{\text{comp}}$, we first use algebraic transformation rules (Table 4) to augment the 25 seed AM-GM problem (Ineq-AMGM) into 100 problems built on the 25 seed problem. Then, we apply composition rules (Table 3) to randomly generate 5000 composed problems.
@@
We observe that if we directly fine-tune Goedel-Prover-SFT or STP on ${\mathcal{D}}_{\text{comp}}$, provers’ performances degrade a lot, possibly because there is a huge distribution shift between ${\mathcal{D}}_{\text{comp}}$ and the models’ training data. In order to mitigate this issue, we mix 5k training data from Goedel-Prover-SFT [Lin et al., 2025] and fine-tune the same model, mitigating the huge discrepancy between ${\mathcal{D}}_{\text{comp}}$ and the training data. The performance on MiniF2F after fine-tuning is retained.
The detailed evaluation results for the fine-tuning experiments is shown in Table 7.
@@
Model Budget Ineq-AMGM Ineq-Cauchy Ineq-MISC Valid Type I Type II Valid Type I Type II Valid Type I Type II Goedel-Prover-SFT [Lin et al., 2025] 32 48.62.9 0.41.2 14.03.2 34.82.5 12.43.5 21.53.4 47.01.7 14.43.1 24.61.9 64 50.62.6 0.81.6 16.62.8 36.21.9 15.83.4 24.62.9 47.80.9 16.62.5 25.51.9 128 52.21.4 1.31.9 18.62.2 37.11.8 19.42.9 26.91.8 48.00.0 17.92.6 26.42.5 3200 60.0 4.0 24.0 40.0 32.0 28.0 48.0 24.0 36.0 Cont. ft on ${\mathcal{D}}_{\text{comp}}$ +5k Goedel SFT Data 32 89.82.7 33.25.3 14.73.6 30.43.1 7.63.7 17.83.4 37.72.5 5.22.4 4.31.8 64 91.51.5 38.72.9 17.43.4 32.62.4 11.24.1 21.13.3 39.51.3 6.52.9 5.22.2 128 92.00.0 42.93.1 19.73.0 34.92.1 15.23.6 23.52.8 40.00.0 8.53.5 6.42.5 3200 92.0 64.0 28.0 44.0 32.0 28.0 40.0 20.0 20.0
@@ Appendix E Prompt Templates
In this section, we document the prompt used for evaluating different models (Section E.1), evaluating by in-context learning (Section E.2), and generating proofs using ICL (Section E.3)
@@ E.1 Prompt template for evaluating different models
E.2 Prompt template for in-context learning experiments (Section 4.1)
@@
E.3 Prompt template for in-context generation (Section 4.2)
@@ Appendix F Example of Model Generations
In this section, we show some of the generations from different models that successfully prove the problem, including Goedel-Prover-SFT [Lin et al., 2025], STP [Dong and Ma, 2025], DeepSeek-R1-Distill-Qwen-32B [Guo et al., 2025], Kimina-Distill-7B [Wang et al., 2025], and DeepSeek-Prover-V2-7B [Ren et al., 2025].
@@ F.5 Example generations from DeepSeek-Prover-V2-7B
DeepSeek-Prover-V2-7B generates code with very similar style to that of DeepSeek-R1-Distill-Qwen-32B, but has much stronger ability. Following is an example from Ineq-Mix that DeepSeek-Prover-V2-7B successfully solves. The problem is constructed by first composing two problems (Table 3) and then applying an algebraic transformation on both the LHS and RHS of the inequality. We can see that DeepSeek-Prover-V2-7B generates very structured code, and decomposes the problem into sub-parts.
@@
7pt
theorem alge_whole_p70 (a b c : ℝ) ( ha : a > 0 ) ( hb : b > 0 ) ( hc : c > 0 ) ( h0 : a + 2 * b = 3 ): ((3:ℝ) * (1) + (2:ℝ) * (( a + b + c ) * ( 1 / a + 1 / b + 1 / c )))^3 ((3:ℝ) * (a * b^2) + (2:ℝ) * (9))^3 := by

### v2

@@ 4 Probing and Mitigating Compositional Challenges
To better understand and potentially mitigate the compositional failures observed in Section˜3, we conduct two ablation studies. First, we explore whether our curated benchmark becomes easier when models’ sizes are scaled up (Section˜4.1), which tests if the compositional failure is unique to relatively small models and can be easily solved by scaling up. Next, we investigate where our benchmark becomes easier when models are provided with the formal proof of the seed problem as part of the input (Section˜4.2). This tests whether models can leverage in-context demonstrations to generalize to the transformed variants. Building on this idea, we then construct synthetic data that explicitly encodes compositional structure and use it to fine-tune the theorem prover. This examines whether the lack of compositional ability can be remedied through supervised fine-tuning alone (Section˜4.3).
4.1 Scaling-up model parameters
In Section˜3, we mostly test theorem provers with sizes at most 32B, and we do not test provers with different sizes within the same family. Thus, it is natural to ask if the compositional weakness pinpointed in Section˜3 can be easily solved by scaling up model parameters.
Table 2: Performance of models with different sizes on Ineq-Comp under pass@32.
Model Seed Type I Type II DeepSeek-Prover-V2-7B (Ren et al., 2025) 66.23 46.97 42.1 DeepSeek-Prover-V2-671B 88.0 68.33 70.67 Goedel-Prover-V2-8B (Lin et al., 2025b) 69.0 35.0 49.0 Goedel-Prover-V2-32B 90.0 68.0 83.0
To answer the question, we test DeepSeek-Prover-V2 7B and its 671B counterpart (Ren et al., 2025), as well as Goedel-Prover-V2 8B and its 32B counterpart (Lin et al., 2025b), under pass@32 on Ineq-Comp. Table˜2 summarizes the result, which shows that although scaling up the model size improves the absolute performance, the relative gap between the Type I, II problems and the seed problems still exists, implying that the compositional weakness in formal theorem proving is not easily solvable by brute-force scaling up model parameters.
@@ 4.2 In-context learning
Table˜3 reports the pass@$N$ accuracy across a range of open-source and proprietary models on Ineq-AMGM, Ineq-Cauchy, and Ineq-MISC problems under different generation budgets. Despite having access to the seed solution, all models continue to struggle with the transformed variants. For example, Qwen2.5-Coder-32B-Instruct achieves only 40.0% accuracy on Type II AM-GM problems and 20.0% on Type II Cauchy–Schwarz problems at 128 attempts. DeepSeek-R1-Distill-Qwen-32B (with thinking) performs similarly, reaching just 16.0% on Type I AM-GM problems. Proprietary models such as GPT-4o and Claude 3.7 Sonnet show mixed results, further illustrating the challenge. These findings suggest that simply providing the seed proof is not enough to enable compositional generalization. Even modest transformations remain difficult, pointing to a fundamental limitation in their ability to transfer and reuse formal reasoning strategies.
Table 3: ICL-based ablation. Performance of various models on Ineq-Simp (Type I and Type II variants) under the in-context learning (ICL) setting. For each test instance, the model is provided with the full formal proof of the corresponding seed problem as part of the input prompt. This setup evaluates whether models can leverage prior solutions to solve transformed variants via compositional generalization. We report pass@NN accuracy (with standard deviation as subscript) across different generation budgets NN. Despite access to a complete solution to the seed problem, all models—including proprietary ones—continue to struggle on both Type I and Type II variants, indicating that the compositional gap cannot be easily closed through in-context demonstrations alone.
@@ 4.3 Fine-tuning
Figure 3: Pipeline for evaluating if composition can be learned through naive fine-tuning. We split the seed problems into ${\mathcal{P}}_{\text{train}}$, which contains 25 problems utilizing AMGM, and ${\mathcal{P}}_{\text{held-out}}$, which contains the remaining 50 problems. The pipeline consists of three steps: (1) generate data by prompting LLM given the proof of seed problems in context and keep only the proofs that pass Lean compilation; (2) fine-tune Goedel-Prover-SFT; (3) evaluate type I and II problems and Ineq-Mix on ${\mathcal{P}}_{\text{held-out}}$.
@@
We split the seed problems into two subsets: 25 AM-GM inequality seed problems as training set and 50 other problems as held-out set. To construct training data, we use our automated framework (Ineq-Mix) to generated 15k problems incorporating both algebraic and compositional transformations base on the AM-GM seed problems. Then we used in-context learning (ICL) to prompt Qwen2.5-Coder-32B-Instruct with seed problems and their verified Lean 4 proofs. We retained approximately 8k compilable outputs and fine-tuned Goedel-Prover-SFT on this data. See Figure˜3 for an overview of the process and Appendix˜D for setup details.
@@ 5 Related Works
ATP has been a long-standing goal in AI (Wu, 2008). Modern theorem provers use tactic-based proof search and premise selection within these systems (e.g. Lean’s tactics in Mathlib environment), but they struggle with enormous search spaces and limited training data for complex theorems (Polu and Sutskever, 2020; Yang et al., 2023). More recently, more theorem provers came out based on large language models (Xin et al., 2024; Lin et al., 2025a; Dong and Ma, 2025; Wang et al., 2025; Ren et al., 2025), and also possibly incorporated tree-search techniques (Wu et al., 2024; Li et al., 2024; Xin et al., 2025).
@@ Appendix B More Details on Ineq-Mix
Positive RHS: $\text{if rhs }g(x_{1},\dots,x_{n})\text{ is guaranteed to be positive.}$
@@
Table 4: (A list of composition methods for two inequality problems.) Given two inequality problems $P_{1},P_{2}$, where $P_{1}$ has variables $X=\{x_{1},\dots,x_{m}\}$, with the statement: given condition on the variables $C_{1}(X)$, the following inequality holds $f_{1}(X)\geq g_{1}(X)$, and $P_{2}$ has variables $X$, with the statement: given condition on the variables $C_{2}(X)$, the following inequality holds $f_{2}(X)\geq g_{2}(X)$, the following list defines several composition operations to generate a new inequality problem. If problems $P_{1}$ and $P_{2}$ have different set of variables, we can always lift the problems and make sure they are defined on the same variable space. Following are the composition methods we considered to combine $P_{1}$ and $P_{2}$ together. If $g_{1}(X)$ and $g_{2}(X)$ are both greater than 0, there are more composition methods. We also make sure that $C_{1}(X)\cap C_{2}(X)$ is non-empty after composition.
@@
Now we show the details of how to compose two inequality problems together or add algebraic transformation to create new inequality problems in Ineq-Mix. Recalled that as mentioned in Section˜2.2, Ineq-Mix consists of transformation/composition rules from three categories: (1) compositional transformations, where (different) inequalities are combined through operations like addition, multiplication, or taking the maximum or minimum of both sides; (2) variable-level algebraic transformations, where we replace variables with other algebraic expressions; and (3) problem-level algebraic transformations, where we apply a monotone function to both sides of an inequality, such as taking exponentials or logarithms. We now introduce them by category.
Compositional transformations Composition transformations denote the rules that given two inequality problems, we compose them together to generate a new inequality, where the two inequality problems serve as subparts for the new inequality problem. Table˜4 shows the full list of composition rules. Note that the rules are classified into two categories to make sure mathematical correctness: the rules require the right-hand side to be positive, and the rules without this requirement. Also note that if the two problems both has positive right-hand side in the statement, then after the compositions defined in Table˜4, the new problem also has a positive right-hand side.
@@
Variable-level algebraic transformations Variable-level algebraic transformations denote the rules that apply algebraic transformations on the variables of a given inequality problem. For example, replacing $x_{i}$ in the original inequality problem into $x_{i}^{2}$, for all $x_{i}\in X$ where $X=\{x_{1},\dots,x_{m}\}$ denotes the variable set. Table˜5 summarizes the variable-level algebraic transformations used for Ineq-Mix. Similar to compositional transformations, there are also two categories for variable-level algebraic transformations. The first category includes the transformations that are “one-to-one” mapping under the basic assumption $X\succ 0$, like taking the square root of all variables, and squaring all variables. Since these transformations are one-to-one, as long as the original problem is feasible, the new problem that applies these rules is also feasible. We also include some transformation rules that are not one-to-one under the basic assumption $X\succ 0$ (but are also elegant transformations). However to ensure the correctness, these rules can only be applied to problems where there is only the basic assumption $X\succ 0$ and no additional assumptions.
@@
Table˜6 summarizes all the problem-level algebraic transformations used in Ineq-Mix.
@@ Appendix C Detailed Model Information
The full list of models used in our experiments, including the version (for API models) and Huggingface-link, is shown in Table˜7.
Table 7: List of models tested in Section˜3 or Section˜4. We use “Kimina-Distill-7B” to denote “Kimina-Prover-Preview-Distill-7B” in the paper.
@@ D.1 Lean 4 version
For evaluation on models except InternLM2.5-StepProver, we stick to the Lean 4 and Mathlib version used in Xin et al. [2024], since Goedel-Prover-SFT [Lin et al., 2025a] and STP [Dong and Ma, 2025] are trained based on DeepSeek-Prover-V1.5-RL [Xin et al., 2024]. Besides, we also nearly recover the reported performance of Kimina-Distill-7B [Wang et al., 2025] and DeepSeek-Prover-V2 [Ren et al., 2025] on MiniF2F benchmark using Lean 4 and Mathlib version in Xin et al. [2024]. We also observe decent performance of DeepSeek-R1-Distill-Qwen-32B on MiniF2F.
When evaluating InternLM2.5-StepProver, we use the Lean 4 version following InternLM2.5-StepProver [Wu et al., 2024] from the Github repository 222 https://github.com/haoyuzhao123/LeanIneqComp-Dojo traceable by LeanDojo [Yang et al., 2023].
@@ D.4 Detailed results for fine-tuning Goedel-Prover-SFT
Please refer to Figure˜3 for an illustration for our fine-tuning experiments pipeline. To generate the fine-tuning data ${\mathcal{D}}_{\text{comp}}$, we first use algebraic transformation rules (Table˜5) to augment the 25 seed AM-GM problem (Ineq-AMGM) into 100 problems built on the 25 seed problem. Then, we apply composition rules (Table˜4) to randomly generate 5000 composed problems.
@@
We observe that if we directly fine-tune Goedel-Prover-SFT or STP on ${\mathcal{D}}_{\text{comp}}$, provers’ performances degrade a lot, possibly because there is a huge distribution shift between ${\mathcal{D}}_{\text{comp}}$ and the models’ training data. In order to mitigate this issue, we mix 5k training data from Goedel-Prover-SFT [Lin et al., 2025a] and fine-tune the same model, mitigating the huge discrepancy between ${\mathcal{D}}_{\text{comp}}$ and the training data. The performance on MiniF2F after fine-tuning is retained.
The detailed evaluation results for the fine-tuning experiments is shown in Table˜8.
@@
Model Budget Ineq-AMGM Ineq-Cauchy Ineq-MISC Valid Type I Type II Valid Type I Type II Valid Type I Type II Goedel-Prover-SFT [Lin et al., 2025a] 32 48.62.9 0.41.2 14.03.2 34.82.5 12.43.5 21.53.4 47.01.7 14.43.1 24.61.9 64 50.62.6 0.81.6 16.62.8 36.21.9 15.83.4 24.62.9 47.80.9 16.62.5 25.51.9 128 52.21.4 1.31.9 18.62.2 37.11.8 19.42.9 26.91.8 48.00.0 17.92.6 26.42.5 3200 60.0 4.0 24.0 40.0 32.0 28.0 48.0 24.0 36.0 Cont. ft on ${\mathcal{D}}_{\text{comp}}$ +5k Goedel SFT Data 32 89.82.7 33.25.3 14.73.6 30.43.1 7.63.7 17.83.4 37.72.5 5.22.4 4.31.8 64 91.51.5 38.72.9 17.43.4 32.62.4 11.24.1 21.13.3 39.51.3 6.52.9 5.22.2 128 92.00.0 42.93.1 19.73.0 34.92.1 15.23.6 23.52.8 40.00.0 8.53.5 6.42.5 3200 92.0 64.0 28.0 44.0 32.0 28.0 40.0 20.0 20.0
@@ Appendix E Prompt Templates
In this section, we document the prompt used for evaluating different models (Section˜E.1), evaluating by in-context learning (Section˜E.2), and generating proofs using ICL (Section˜E.3)
@@ E.1 Prompt template for evaluating different models
E.2 Prompt template for in-context learning experiments (Section˜4.2)
@@
E.3 Prompt template for in-context generation (Section˜4.3)
@@ Appendix F Example of Model Generations
In this section, we show some of the generations from different models that successfully prove the problem, including Goedel-Prover-SFT [Lin et al., 2025a], STP [Dong and Ma, 2025], DeepSeek-R1-Distill-Qwen-32B [Guo et al., 2025], Kimina-Distill-7B [Wang et al., 2025], and DeepSeek-Prover-V2-7B [Ren et al., 2025].
@@ F.5 Example generations from DeepSeek-Prover-V2-7B
DeepSeek-Prover-V2-7B generates code with very similar style to that of DeepSeek-R1-Distill-Qwen-32B, but has much stronger ability. Following is an example from Ineq-Mix that DeepSeek-Prover-V2-7B successfully solves. The problem is constructed by first composing two problems (Table˜4) and then applying an algebraic transformation on both the LHS and RHS of the inequality. We can see that DeepSeek-Prover-V2-7B generates very structured code, and decomposes the problem into sub-parts.
@@
theorem alge_whole_p70 (a b c : ℝ) ( ha : a > 0 ) ( hb : b > 0 ) ( hc : c > 0 ) ( h0 : a + 2 * b = 3 ): ((3:ℝ) * (1) + (2:ℝ) * $($ a + b + c ) * ( 1 / a + 1 / b + 1 / c )))^3 ((3:ℝ) * (a * b^2) + (2:ℝ) * (9))^3 := by